# Phase Transition and Nuclear Symmetry Energy from Neutron Star Observations – In Light of PSR J0614–3329

### v1

@@ 4 Summary
Thanks to an expanding array of multi-messenger observations, our understanding of the neutron star EOS continues to improve. The NICER mission has now provided radius measurements for four pulsars, and although each individual measurement still has substantial uncertainty, collectively these data have significantly reduced the allowed EOS parameter space. In this work, we revisited the question of a possible phase transition inside the NSs, using flexible EOS models and incorporating the latest observational inputs (including the new NICER result for PSR J0614–3329). We extend our previous study (Han et al., 2023) by considering an explicit FOPT in a model-agnostic way. Comparing EOS ensembles with and without an explicit PT yields a Bayes factor of 2.3, marginally favoring the PT hypothesis. While the result is generally consistent with the conclusion of Han et al. (2023) that $c_{s}^{2}$ approaches zero near the central density of the most massive NS, the inferred PT points suggest two scenarios: low-density PT at densities below $2\,n_{s}$ and more prominently, high-density PT at densities above $\sim 4\,n_{s}$. We further study the critical mass at which the PT occurs at the center of NSs. In the low‐density scenario, the transition onset corresponds to masses of $0.44^{+0.29}_{-0.22}\,M_{\odot}$, with the upper bound falling below the precisely measured mass of the lightest observed neutron star (the companion to PSR J0453+1559 at $1.174\pm 0.04\,M_{\odot}$ (Martinez et al., 2015)). In the high‐density scenario, the PT nearly coincides with the maximum‐mass configuration, with a mass of $2.24^{+0.08}_{-0.07}M_{\odot}$ ($M_{\rm TOV}=2.25^{+0.08}_{-0.07}M_{\odot}$) and a radius of $11.65^{+0.56}_{-0.42}$ km ($R_{\rm TOV}=11.59^{+0.57}_{-0.46}$ km). Extremely large density jumps are ruled out by the pQCD constraints. Furthermore, we find that $n_{\rm PT}$ is well-correlated with $R_{\rm TOV}$. This suggests that a precise determination of $R_{\rm TOV}$ in the future (for example, via post-merger gravitational-wave signatures or other novel methods) could provide critical insight into the presence of a phase transition in NS matter. We have also obtained tighter constraints on the symmetry energy parameters. The slope parameter $L$ is constrained to be $40.18^{+19.25}_{-14.30}$ MeV in our analysis. Interestingly, we observe that $L$ correlates with the difference in radii between $1.4\,M_{\odot}$ and $2.0\,M_{\odot}$ NSs. This correlation means that improved radius measurements for pulsars of different masses (e.g., continued observations of PSR J0437–4715 at $\sim 1.4\,M_{\odot}$ and a future NICER or eXTP measurement of a heavy pulsar around $2\,M_{\odot}$) could further tighten the constraints on $L$. Such astrophysical measurements of $L$ would be especially valuable given the current tension between the PREX-II neutron-skin result and other nuclear experimental results. Finally, we examined the effect of the new PSR J0614–3329 data on the inferred EOS. We found that including this latest NICER measurement shifts the preferred mass-radius relation toward smaller radii with average $\sim 0.3$ km, and slightly softens the pressure at intermediate densities relative to our previous results. Many bulk NS properties (such as $R_{\rm TOV}$, $R_{1.4}$, and $R_{2.0}$) are now determined with uncertainties on the order of only a few hundred meters in $68\%$ confidence interval. Importantly, the inclusion of the new data and the explicit phase-transition modeling does not significantly change our previously obtained $M_{\rm TOV}$, underscoring the robustness of the maximum-mass constraint near $2.2$-$2.3\,M_{\odot}$ (Fan et al., 2013, 2024; Tang et al., 2025).
@@
Given the results above, the possible observable effect for a strong FOPT (e.g., a sharp radius reduction) is favored to lie outside the mass-radius range accessible to NS observations. Current multi-messenger constraints suggest that it is hard to directly confirm the FOPT from NS observations alone, even with precise mass-radius measurements. Future confirmation of FOPT is expected to be achieved by incorporating both NS observations for the cold EOS and its finite-temperature properties constrained by nuclear experiments. For this purpose, we report the chemical potential for the PT, $\mu_{\rm PT}$, at zero temperature, which also shows two peaks, located at $\sim$ 1000 and $\sim$ 1500 MeV.
@@
Looking ahead, as multi-disciplinary data accumulates from both space telescopes for NS observations (like the continuing operation of NICER and next-generation missions under development, e.g., eXTP (Zhang et al., 2019; Watts et al., 2019; Li et al., 2025a)) and terrestrial nuclear experiments, such as the Relativistic Heavy Ion Collider (RHIC), the Large Hadron Collider (LHC), the Super Proton Synchrotron (SPS), the Facility for Antiproton and Ion Research (FAIR), the Nuclotron-based Ion Collider Facility (NICA), and high-intensity heavy-ion accelerator facilities (HIAF) (Bzdak et al., 2020; Zhou et al., 2022; Sorensen et al., 2024). It is promising to resolve the remaining questions regarding possible phase transitions at both zero and finite temperature. These advances would be connected by numerical simulations for binary NS mergers (Bauswein et al., 2019; Most et al., 2019; Weih et al., 2020; Prakash et al., 2024; Huang et al., 2022; Hensh et al., 2024), potentially shedding light on the location of the conjectured critical end point (Ecker et al., 2025; Constantinou et al., 2025).
@@
This work is supported by the National Natural Science Foundation of China under Grants of No. 12233011 and No. 12303056, the Project for Young Scientists in Basic Research (No. YSBR-088) of the Chinese Academy of Sciences, and the Postdoctoral Fellowship Program (GZC20241915) of the China Postdoctoral Science Foundation.

### v2

@@ IV Summary
Thanks to an expanding array of multimessenger observations, our understanding of the neutron star EOS continues to improve. The NICER mission has now provided radius measurements for four pulsars, and although each individual measurement still has substantial uncertainty, collectively these data have significantly reduced the allowed EOS parameter space. In this work, we revisited the question of a possible phase transition inside the NSs, using flexible EOS models and incorporating the latest observational inputs (including the new NICER result for PSR J0614–3329). We extend our previous study (Han et al., 2023) by considering an explicit FOPT in a model-agnostic way. Comparing EOS ensembles with and without an explicit PT yields a Bayes factor of 2.3, marginally favoring the PT hypothesis. While the result is generally consistent with the conclusion of Ref. (Han et al., 2023) that $c_{s}^{2}$ approaches zero near the central density of the most massive NS, the inferred PT points suggest two scenarios: low-density PT at densities below $2\,n_{s}$ and more prominently, high-density PT at densities above $\sim 4\,n_{s}$. We further study the critical mass at which the PT occurs at the center of NSs. In the low‐density scenario, the transition onset corresponds to masses of $0.44^{+0.29}_{-0.22}\,M_{\odot}$, with the upper bound falling below the precisely measured mass of the lightest observed neutron star (the companion to PSR J0453+1559 at $1.174\pm 0.04\,M_{\odot}$ (Martinez et al., 2015)). In the high‐density scenario, the PT nearly coincides with the maximum‐mass configuration, with a mass of $2.24^{+0.08}_{-0.07}M_{\odot}$ ($M_{\rm TOV}=2.25^{+0.08}_{-0.07}M_{\odot}$) and a radius of $11.65^{+0.56}_{-0.42}$ km ($R_{\rm TOV}=11.59^{+0.57}_{-0.46}$ km). Extremely large density jumps are ruled out by the pQCD constraints. Furthermore, we find that $n_{\rm PT}$ is well correlated with $R_{\rm TOV}$. This suggests that a precise determination of $R_{\rm TOV}$ in the future (for example, via postmerger gravitational-wave signatures or other novel methods) could provide critical insight into the presence of a phase transition in NS matter. We have also obtained tighter constraints on the symmetry energy parameters. The slope parameter $L$ is constrained to be $40.18^{+19.25}_{-14.30}$ MeV in our analysis. Interestingly, we observe that $L$ correlates with the difference in radii between $1.4\,M_{\odot}$ and $2.0\,M_{\odot}$ NSs. This correlation means that improved radius measurements for pulsars of different masses (e.g., continued observations of PSR J0437–4715 at $\sim 1.4\,M_{\odot}$ and a future NICER or enhanced X-ray Timing and Polarimetry (eXTP) measurement of a heavy pulsar around $2\,M_{\odot}$) could further tighten the constraints on $L$. Such astrophysical measurements of $L$ would be especially valuable given the current tension between the PREX-II neutron-skin result and other nuclear experimental results. Finally, we examined the effect of the new PSR J0614–3329 data on the inferred EOS. We found that including this latest NICER measurement shifts the preferred mass-radius relation toward smaller radii by an average $\sim 0.3$ km, and slightly softens the pressure at intermediate densities relative to our previous results. Many bulk NS properties (such as $R_{\rm TOV}$, $R_{1.4}$, and $R_{2.0}$) are now determined with uncertainties on the order of only a few hundred meters in the $68\%$ confidence interval. Importantly, the inclusion of the new data and the explicit phase-transition modeling does not significantly change our previously obtained $M_{\rm TOV}$, underscoring the robustness of the maximum-mass constraint near $2.2$-$2.3\,M_{\odot}$ (Fan et al., 2013, 2024; Tang et al., 2025).
@@
Given the results above, the possible observable effect for a strong FOPT (e.g., a sharp radius reduction) is favored to lie outside the mass-radius range accessible to NS observations. Current multimessenger constraints suggest that it is hard to directly confirm the FOPT from NS observations alone, even with precise mass-radius measurements. Future confirmation of FOPT is expected to be achieved by incorporating both NS observations for the cold EOS and its finite-temperature properties constrained by nuclear experiments. For this purpose, we report the chemical potential for the PT, $\mu_{\rm PT}$, at zero temperature, which also shows two peaks, located at $\sim$ 1000 and $\sim$ 1500 MeV.
@@
Looking ahead, as multidisciplinary data accumulate from both space telescopes for NS observations (like the continuing operation of NICER and next-generation missions under development, e.g., eXTP (Zhang et al., 2019; Watts et al., 2019; Li et al., 2025d)) and terrestrial nuclear experiments, such as the Relativistic Heavy Ion Collider, the Large Hadron Collider, the Super Proton Synchrotron, the Facility for Antiproton and Ion Research, the Nuclotron-based Ion Collider Facility, and high-intensity heavy-ion accelerator facilities (Bzdak et al., 2020; Zhou et al., 2022; Sorensen et al., 2024). It is promising to resolve the remaining questions regarding possible phase transitions at both zero and finite temperature. These advances would be connected by numerical simulations for binary NS mergers (Bauswein et al., 2019; Most et al., 2019; Weih et al., 2020; Prakash et al., 2024; Huang et al., 2022; Hensh et al., 2024), potentially shedding light on the location of the conjectured critical end point (Ecker et al., 2025; Constantinou et al., 2025).
Acknowledgements.
@@
This work is supported by the National Natural Science Foundation of China under Grants No. 12233011 and No. 12303056, the Project for Young Scientists in Basic Research (No. YSBR-088) of the Chinese Academy of Sciences, and the Postdoctoral Fellowship Program (No. GZC20241915) of the China Postdoctoral Science Foundation.
DATA AVAILABILITY
The data that support the findings of this article are openly available (https://doi.org/10.5281/zenodo.17149145).